a flowerpot falls from a windowsill 25.0 m above the sidewalk. How fast is the flowerpot moving when it strikes the ground?
Answer #1
x(t) = 1/2*a*t^2

-25 = -4.9t^2

t = 2.2587

v = at + v0

v(2.2587) = -9.8*2.2587 = -22.135 m/s <ANSWER

PLEASE RATE.
answered by: Rosh